The length of a rectangle is 3 times the width. The perimeter is 96 cm. Find the width and length.
can you please include the formula

Answer by tw23279(14)

So, the perimeter for a rectangle = 2(length + width)
And we know that the length is 3 times the width.
So length, l = 3w.
Substitute.
We get that 2(3w + w) = 96 (And the 96 is the 96 cm provided by the problem)
Distribute!
6w + 2w = 96
8w = 96
w = 12 cm
And so recall that length = 3w
length = 3(12)
length = 36 cm